OpenStudy (anonymous):

can someone show me the steps of solving this problem: 5x-2y=14 -x-3y=4 ...solve by the method of elimination

OpenStudy (anonymous):

multiply the second equation all the way across by 5, then add it to the first equation the \(x\) term will be eliminated

OpenStudy (anonymous):

can you please show me how to do that

OpenStudy (callisto):

5x-2y=14 -(1) -x-3y=4 -(2) (2)x5 +(1) 5(-x-3y) + (5x-2y) = 5(4) + 14 <- solve this and you can get the value of y.

OpenStudy (anonymous):

can you show we how to solve it